OpenStudy (anonymous):

can someone help me please??? solve the system using an equivalent system. -2x - y = 0 3x - y= 10 (4,-2) (-2,4) (2,-4) (-4,2)

OpenStudy (anonymous):

If you have -2x-y=0 first you need to get y by itself so add it to the other side. then you have -2x=y. you can then plug this into the other equation once you get y by itself there too. so you would end up with y=-10+3x -2x=-10+3x 10=5x x=2 y=-4 put x back into the first equation -2x=y the answer is (2,-4)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no problem